# Cosmic filament spin - I: a comparative study in observation

### v1

@@ 1 Introduction
A variety of methods have been developed to identify and classify cosmic structures such as filaments, clusters, and large-scale features in the cosmic web. These methods use diverse approaches, including graph-based algorithms (e.g., the Minimal Spanning Tree method, Alpaslan et al. 2014a), stochastic modeling (e.g., Bisous, Tempel et al. 2014, 2016; FINE, González & Padilla 2010), and tensor-based Hessian analysis (e.g., T-web, Forero-Romero et al. 2009; V-web, Hoffman et al. 2012; CLASSIC, Kitaura & Angulo 2012). Advanced scale-space techniques are utilized by NEXUS+ (Cautun et al. 2013) and MMF-2 (Aragón-Calvo et al. 2007; Aragon-Calvo & Yang 2014), while topological approaches are adopted by Spineweb (Aragón-Calvo et al. 2010b) and DisPerSE (Sousbie 2011). Phase-space analyses are used in ORIGAMI (Falck et al. 2012; Falck & Neyrinck 2015) and MSWA (Ramachandra & Shandarin 2015). These methods vary in the types of structures they identify, such as filaments, clusters, or the entire cosmic web, and in their sensitivity to different datasets, including particles and galaxy/haloes. This variety highlights ongoing efforts to improve cosmic structure detection and advance our understanding of large-scale structure formation. However, the complexity of filamentary structures results in inconsistencies among the outcomes of different algorithms. Currently, there is no universally accepted method for defining filaments. For a detailed review of these classification techniques, we refer readers to (Libeskind et al., 2018).
@@ 2 Data and Methodology
Table 1: Overview of the galaxy and filament catalogues used in this study. Each column, from left to right, contains the referenced sample name, tracer galaxy, filament finder, number of filaments, galaxy employed to determine filament spin (measure galaxy), and the number of galaxies. In ‘Sample-3’ and ‘Sample-3’, the redshifts of galaxies in Measure Galaxy (Alam et al., 2015) are not corrected for redshift space distortion (RSD).
@@
Sample-3 Zandivarez & Martínez (2011) M16 (Martínez et al., 2015) 3,094 Alam et al. (2015), RSD
Sample-4 Alam et al. (2015) DisPerSE (Sousbie, 2011) 26,974 Tempel et al. (2017a) 584,449
@@ 2.1.2 Sample-2
The second catalogue of filaments, referred to as ‘Sample-2’, used in this study is derived from (Pereyra et al., 2020). It is constructed using a cosmological filament identification algorithm, which employs a technique from graph theory known as the minimal spanning tree (MST, Barrow et al., 1985). The MST is an algorithm in graph theory that is used to find the spanning tree with the smallest total weight in a weighted graph. Based on previous investigations (Barrow et al., 1985; Colberg, 2007; Alpaslan et al., 2014b), Pereyra et al. (2020) adopted the same definitions for identifying filaments.
@@ 2.1.3 Sample-3
In ‘Sample-3’, filaments are identified from a galaxy group catalogue (Zandivarez & Martínez, 2011). The filament catalogue (Martínez et al., 2015) was created considering pairs of groups, including only those with virial masses larger than $\rm M_{vir}>10^{13.5}$, within the redshift interval $0.150.05<z<$ 0.15. Filaments are represented as straight lines connecting nearby galaxy groups, thus inherently ignoring any curved or branched formations. Three selection criteria are applied: A pair of groups is connected by a filament if:
@@
Clearly, based on the definition of filaments, this filament sample is incomplete. Nevertheless, (Martínez et al., 2015) aim to investigate the impact of filaments on galaxies that are falling into groups. Following the identification process, a catalogue was compiled, listing 3,094 filaments along with their physical characteristics, including the distance between groups (filament length), the count of member galaxies, and more, for subsequent studies.
@@ 2.2 The Measure Galaxy samples
To address how the filament spin signal is influenced by galaxy samples, we used three galaxy samples for spin signal measurements, as listed in the last two columns of Table 1. The first galaxy sample, which was only used in the ‘Sample-1’ comes from Tempel et al. (2012), contains 576,493 galaxies. The galaxy sample used in ‘Sample-2’ and ‘Sample-3’ is the same, constructed from Alam et al. (2015), contains 584,434 galaxies. The third galaxy sample (Tempel et al., 2017a), used for ‘Sample-4’ and identical to the galaxy sample in Wang et al. (2021), was derived from a publicly accessible catalogue (Tempel et al., 2017a) and was compiled from (Eisenstein et al., 2011; Alam et al., 2015), contains 584,449 galaxies.
@@
In our study, the redshifts of galaxies in ‘Sample-2’ and ‘Sample-3’ are uncorrected for RSD, whereas those in ‘Sample-1’ and ‘Sample-4’ are corrected. To mitigate the impact of RSD, Tempel et al. (2017b) provides an effective approach. This approach involves associating galaxies from the group catalogue with groups identified using a modified Friends-of-Friends (FoF) halo finder (Huchra & Geller, 1982; Davis et al., 1985; Springel et al., 2001; Dolag et al., 2009). Subsequently, corrections are applied to galaxy positions along the line of sight, taking into account both the radial velocity dispersion of the group and its angular size on the sky plane.
@@ 4 Summary and Discussion
Here, as the first step of the Cosmic Filament Spin project , we have examined filament signals in redshift space. We acknowledge that asymmetrical redshift-space filament shapes not completely arising from spin (present, e.g. even in real space) might contribute in some cases to the observed spin signal. A detailed investigation of this influence, including an evaluation of possible false positives, will be addressed in a forthcoming paper. Also, much further study is necessary of the origin and evolution of filament spin, and how consistent the observations are with existing theoretical ideas.

### v3

@@ 1 Introduction
A variety of methods have been developed to identify and classify cosmic structures such as filaments, clusters, and large-scale features in the cosmic web. These methods use diverse approaches, including graph-based algorithms (e.g., the Minimal Spanning Tree method, Alpaslan et al. 2014), stochastic modeling (e.g., Bisous, Tempel et al. 2014, 2016; FINE, González & Padilla 2010), and tensor-based Hessian analysis (e.g., T-web, Forero-Romero et al. 2009; V-web, Hoffman et al. 2012; CLASSIC, Kitaura & Angulo 2012). Advanced scale-space techniques are utilized by NEXUS+ (Cautun et al. 2013) and MMF-2 (Aragón-Calvo et al. 2007; Aragon-Calvo & Yang 2014), while topological approaches are adopted by Spineweb (Aragón-Calvo et al. 2010b) and DisPerSE (Sousbie 2011). Phase-space analyses are used in ORIGAMI (Falck et al. 2012; Falck & Neyrinck 2015) and MSWA (Ramachandra & Shandarin 2015). These methods vary in the types of structures they identify, such as filaments, clusters, or the entire cosmic web, and in their sensitivity to different datasets, including particles and galaxy/haloes. This variety highlights ongoing efforts to improve cosmic structure detection and advance our understanding of large-scale structure formation. However, the complexity of filamentary structures results in inconsistencies among the outcomes of different algorithms. Currently, there is no universally accepted method for defining filaments. For a detailed review of these classification techniques, we refer readers to (Libeskind et al., 2018).
@@ 2 Data and Methodology
Table 1: Overview of the galaxy and filament catalogues used in this study. Each column, from left to right, contains the referenced sample name, tracer galaxy, filament finder, number of filaments, galaxy employed to determine filament spin (measure galaxy), and the number of galaxies. In ‘Sample-2’ and ‘Sample-3’, the redshifts of galaxies in Measure Galaxy (Alam et al., 2015) are not corrected for redshift space distortion (RSD).
@@
Sample-3 Zandivarez & Martínez (2011) M16 (Martínez et al., 2016) 3,094 Alam et al. (2015), RSD
Sample-4 Alam et al. (2015) DisPerSE (Sousbie, 2011) 26,974 Tempel et al. (2017) 584,449
@@ 2.1.2 Sample-2
The second catalogue of filaments, referred to as ‘Sample-2’, used in this study is derived from (Pereyra et al., 2020). It is constructed using a cosmological filament identification algorithm, which employs a technique from graph theory known as the minimal spanning tree (MST, Barrow et al., 1985). The MST is an algorithm in graph theory that is used to find the spanning tree with the smallest total weight in a weighted graph. Based on previous investigations (Barrow et al., 1985; Colberg, 2007; Alpaslan et al., 2014), Pereyra et al. (2020) adopted the same definitions for identifying filaments.
@@ 2.1.3 Sample-3
In ‘Sample-3’, filaments are identified from a galaxy group catalogue (Zandivarez & Martínez, 2011). The filament catalogue (Martínez et al., 2016) was created considering pairs of groups, including only those with virial masses larger than $\rm M_{vir}>10^{13.5}$, within the redshift interval $0.150.05<z<$ 0.15. Filaments are represented as straight lines connecting nearby galaxy groups, thus inherently ignoring any curved or branched formations. Three selection criteria are applied: A pair of groups is connected by a filament if:
@@
Clearly, based on the definition of filaments, this filament sample is incomplete. Nevertheless, (Martínez et al., 2016) aim to investigate the impact of filaments on galaxies that are falling into groups. Following the identification process, a catalogue was compiled, listing 3,094 filaments along with their physical characteristics, including the distance between groups (filament length), the count of member galaxies, and more, for subsequent studies.
@@ 2.2 The Measure Galaxy samples
To address how the filament spin signal is influenced by galaxy samples, we used three galaxy samples for spin signal measurements, as listed in the last two columns of Table 1. The first galaxy sample, which was only used in the ‘Sample-1’ comes from Tempel et al. (2012), contains 576,493 galaxies. The galaxy sample used in ‘Sample-2’ and ‘Sample-3’ is the same, constructed from Alam et al. (2015), contains 584,434 galaxies. The third galaxy sample (Tempel et al., 2017), used for ‘Sample-4’ and identical to the galaxy sample in Wang et al. (2021), was derived from a publicly accessible catalogue (Tempel et al., 2017) and was compiled from (Eisenstein et al., 2011; Alam et al., 2015), contains 584,449 galaxies.
@@
In our study, the redshifts of galaxies in ‘Sample-2’ and ‘Sample-3’ are uncorrected for RSD, whereas those in ‘Sample-1’ and ‘Sample-4’ are corrected. To mitigate the impact of RSD, Tempel et al. (2017) provides an effective approach. This approach involves associating galaxies from the group catalogue with groups identified using a modified Friends-of-Friends (FoF) halo finder (Huchra & Geller, 1982; Davis et al., 1985; Springel et al., 2001; Dolag et al., 2009). Subsequently, corrections are applied to galaxy positions along the line of sight, taking into account both the radial velocity dispersion of the group and its angular size on the sky plane.
@@ 4 Summary and Discussion
Here, as the first step of the Cosmic Filament Spin project, we have examined filament signals in redshift space. We acknowledge that asymmetrical redshift-space filament shapes not completely arising from spin (present, e.g. even in real space) might contribute in some cases to the observed spin signal. A detailed investigation of this influence, including an evaluation of possible false positives, will be addressed in a forthcoming paper (Wang et al., 2025). Also, much further study is necessary of the origin and evolution of filament spin, and how consistent the observations are with existing theoretical ideas.